Given that A is obtuse and B is reflex angles and sin A=12/13 and cosB=3/5 without using calculators, evaluate tan(A+B)


1
Expert's answer
2021-09-13T16:37:39-0400

"Sin\space A=\frac{12}{13}"

A is obtuse, A lie in the 2nd quadrant



"Tan\space A=-\frac{12}{5}"

B is a reflex angle (B lie in 3rd or 4thquadrant

"Cos\space B=\frac{3}{5}"



Cos B is positive hence B lie in 4th quadrant.

"Tan\space B=-\frac{4}{3}"

"Tin\space( A+B)=\frac{Tan\space A Tan \space B}{1-Tan \space ATan \space B}=\frac{\frac{-12}{5}\frac{-4}{3}}{1-(\frac{-12}{5})(\frac{-4}{3})}"

"=\frac{\frac{-56}{15}}{\frac{-33}{15}}\\=\frac{56}{33}"